OpenStudy (anonymous):

@hoblos : The equation in point-slope form of the line which passes through (-2, 3) and is parallel to 3x + 2y = 10 is

OpenStudy (hoblos):

3x + 2y = 10 then 2y = -3x +10 then y = -3/2x +5 so the slope is -3/2 applying the formula y-y1 = m(x-x1) y-3 = (-3/2)(x+2)